# Local Advice and Local Decompression

### v1

@@ 2. Related work
In (Fraigniaud et al., 2009), the authors investigate the problem of distributedly $3$-coloring a cycle. Without advice, it is known that this problem requires $\Omega(\log^{*}n)$ rounds (Linial, 1992), and the paper seeks to understand whether advice can help to break this barrier. In the context studied in (Fraigniaud et al., 2009), each node receives some advice as input, and the total advice is measured as the sum of the lengths of the bit-strings given to all nodes. The authors show that, for any constant $k$, $O(n/\log^{(k)}n)$ bits of total advice are not sufficient to beat the $\Omega(\log^{*}n)$ lower bound, where $\log^{(k)}n$ denotes $k$ recursive iterations of $\log n$.
@@
The authors of (Emek et al., 2009) introduced the notion of advice in the context of online algorithms. The authors studied the impact of advice in the competitive ratio of some classical online problems, namely metrical task systems and $k$-server, and they show tradeoffs between the number of bits of advice and the achievable competitive ratio. We refer to the survey (Boyar et al., 2017) for more work on online algorithms with advice.
In (Christiansen et al., 2023), it has been shown that distributed advice can help to design faster dynamic algorithms for coloring. More in detail, the authors design a dynamic data structure that maintains a coloring of a graph, where the number of colors depends on the arboricity of the graph. To do that, authors simulate algorithms that are designed for the distributed setting. To simulate them efficiently in the centralized setting, they consider distributed algorithms with advice, with a type of advice that can be maintained efficiently in the centralized setting.
3. Preliminaries
3.1. Generic Notions
@@
Let $G=(V,E)$ be a graph, where $V$ is the set of nodes and $E$ be the set of edges. We may use the notation $V(G)$ to denote the set of nodes of $G$, and $E(G)$ to denote the set of edges of $G$. With $\operatorname{dist}_{G}(u,v)$ we denote the distance between $u\in V$ and $v\in V$ in $G$, that is, the length of the shortest path between $u$ and $v$ in $G$, where the length of a path is the number of edges of the path. For an integer $k$, the power graph $G^{k}$ of a graph $G$ is the graph that contains the same nodes as $G$, and there is an edge $\{u,v\}$ in $G^{k}$ if and only if $1\leq\operatorname{dist}_{G}(u,v)\leq k$. Two nodes $u$ and $v$ are neighbors in $G$ if there is an edge $\{u,v\}\in E(G)$.
@@
Lemma 0 (Lovász Local Lemma (Shearer, 1985; Spencer, 1977; Erdős and Lovász, 1975)).
@@ 3.2. LOCAL Model
3.3. Locally Checkable Labelings
We formally define an LCL problem $\Pi$ as a tuple $(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},C,r)$ where each element of the tuple is defined as follows. The parameters $\Sigma_{\mathrm{in}}$ and $\Sigma_{\mathrm{out}}$ are finite sets of input and output labels, respectively. The parameter $r$ is a positive integer called checkability radius of $\Pi$, i.e., it determines how far in the graph each node needs to check in order to verify the validity of a given solution. The parameter $C$ determines the constraints of $\Pi$, that is, $C$ is a finite set containing pairs $(v,H)$ where: $H$ is a labeled graph and $v$ is a node of $H$; the eccentricity of $v$ in $H$ is at most $r$; each pair $(v,e)\in V_{H}\times E_{H}$ satisfying that $v$ is an endpoint of $e$ has a label $\ell_{\mathrm{in}}\in\Sigma_{\mathrm{in}}$ and a label $\ell_{\mathrm{out}}\in\Sigma_{\mathrm{out}}$.
Let $\Pi=(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},C,r)$ be an LCL problem and let $G=(V,E)$ be a graph where each node-edge pair $(v,e)$ (where $v$ is an endpoint of $e$) is labeled with a label from $\Sigma_{\mathrm{in}}$. The task of solving $\Pi$ on $G$ requires labeling each node-edge pair $(v,e)\in V\times E$ (where $v$ is an endpoint of $e$) with a label in $\Sigma_{\mathrm{out}}$ such that, for each node $v\in V$ it holds that the graph induced by the nodes at distance at most $r$ from $v$ is isomorphic to some (labeled) graph in $C$.
3.4. Advice Schema
@@
Definition 0 (Advice Schema).
@@
  • There exists an algorithm $\mathcal{A}$ that, for each $G\in\mathcal{G}$, if we label each node $v\in V$ with the bit-string $\ell_{G}(v)$, then $\mathcal{A}$ runs in at most $T$ rounds and outputs a valid solution for $\Pi$, where $T$ is a function of $\Delta$.
@@
  2. (2)
    If a subset of nodes receive bit-strings of the same length, and all the others receive bit-strings of length 00, then the schema is called subset fixed-length.
  3. (3)
    If a subset of nodes receive a bit-string of possibly different positive lengths, and all the others receive bit-strings of length 00, then the schema is called variable-length.
@@
For uniform fixed-length advice schemas where all nodes receive one bit, we define the notion of sparsity, which captures the fact that the ratio between nodes receiving a $1$ and nodes receiving a 00 can be made arbitrarily small.
Definition 0 (Sparse schema).
@@
  • For any graph $G\in\mathcal{G}$, let $n_{0}$ be the amount of nodes to which the schema (i.e., the function $f$ of Definition 2) assigns a 00, and let $n_{1}$ be the amount of nodes to which $f$ assigns a $1$. Then, $\frac{n_{1}}{n_{0}+n_{1}}\leq\varepsilon$.
@@
3.5. Composability
The main idea that we will use to devise our advice schemas is the following. Given some problem $\Pi$, it may be cumbersome to directly define an advice schema for it. Instead, it may be easier to do this operation gradually. More in detail, many problems can be solved as follows: first, find a solution for some subproblem; then, use the solution for the subproblem in order to solve the problem more easily. As a running example, consider the following problem $\Pi$.
@@ 3.5. Composability
  • It is required to output a coloring of the edges, say red and blue, such that each node has the same amount of red and blue incident edges.
@@
  • Let $\Pi_{\mathrm{v}}$ be the problem of computing a $2$-coloring of the nodes, say with colors black and white.
  • Let $\Pi_{\mathrm{e}}$ be the problem of outputting a $2$-coloring of the edges, such that each node has the same amount of red and blue incident edges, assuming that we are given as input a $2$-coloring of the nodes, and assuming that all nodes have even degree.
  • Let $\Pi_{\mathrm{o}}$ be the problem of orienting all edges such that each node has the same amount of incoming and outgoing edges, assuming that all nodes have even degree.
Consider the following algorithm. First, solve $\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$. Then, solve $\Pi_{\mathrm{e}}$, by coloring red the edges oriented from black to white, and by coloring blue the edges oriented from white to black. Observe that this algorithms solves $\Pi$.
In other words, we decomposed $\Pi$ into three different subproblems: two of them are “hard” problems ($\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$), for which some advice is needed if we want to solve them fast, and one ($\Pi_{\mathrm{e}}$) is trivial once we are given as input the solution for the other two. The idea now is devise two different advice schemas for $\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$ separately.
  • For $\Pi_{\mathrm{v}}$ there is a trivial advice schema: just use $1$ bit to encode the $2$-coloring of the nodes.
  • For $\Pi_{\mathrm{o}}$ the advice schema is non-trivial, and we will see how it can be done in Section 5 by using $1$ bit per node.
@@
Once we have a composable schema for all subproblems that are used to solve our problem of interest, we apply, as a black box, a lemma (see Lemma 1) to obtain a single variable-length schema for our problem. Then, again as a black box, we convert such a schema into a uniform fixed-length schema that uses a single bit per node (see Lemma 2).
@@
  • We prove that such properties imply that we can compose many variable-length schemas into a single one that satisfies the same properties (see Lemma 1);
  • We prove that the resulting variable-length schema implies a uniform fixed-length schema that uses a single bit per node (see Lemma 2).
Note that the last two steps are done in a problem-independent way. The conditions that allow to combine multiple variable-length schemas are expressed in Definition 4, that, on a high level, states the following. We want our variable-length schema to be tunable as a function of three parameters $\alpha$, $\gamma$, and $c$. The schema needs to satisfy that in each $\alpha$-radius neighborhood there are at most $\gamma$ bit-holding nodes, and that the amount of bits held by these nodes is upper bounded by $c\alpha/\gamma^{3}$. Hence, a composable schema is a collection of schemas, one for each choice of parameters $\alpha$, $\gamma$, and $c$.
Definition 0 (Composability).
@@
4. LCLs on Graphs with Subexponential Growth

### v2

@@ 2 Additional related work
Fraigniaud, Gavoille, Ilcinkas, and Pelc [22] investigate the problem of distributedly $3$-coloring a cycle. Without advice, it is known that this problem requires $\Omega(\log^{*}n)$ rounds [45], and the paper seeks to understand whether advice can help to break this barrier. In the context studied in [22], each node receives some advice as input, and the total advice is measured as the sum of the lengths of the bit-strings given to all nodes. The authors show that, for any constant $k$, $O(n/\log^{(k)}n)$ bits of total advice are not sufficient to beat the $\Omega(\log^{*}n)$ lower bound, where $\log^{(k)}n$ denotes $k$ recursive iterations of $\log n$.
Algorithm design.
Parnas and Ron [54] show that one can use distributed algorithms to derive (by simulation) Local Computation Algorithms. Quite naturally, one can use the same approach to derive parallel graph algorithms or graph algorithms for dynamic data sets (or graph algorithms for any model of computation that can leverage the fact that the output is defined by some small neighborhood of a vertex or edge). However, as is, the reduction from [54] did not immediately imply many new state-of-the-art algorithms, as the size of the neighborhood grows exponentially with locality, and, for many problems, the locality of the fastest-known algorithms is fairly large. One of the ways to address this issue is to rely on analysis of algorithms with advice.
Christiansen, Nowicki, and Rotenberg [12] use distributed algorithms with advice to design faster dynamic algorithms for vertex coloring, by combining dynamic graph orientations with ideas for simulating distributed algorithms on the resulting directed graph.
While this approach in principle can be used in the design of parallel algorithms, so far it is used implicitly or in a somewhat trivial way (e.g. Ghaffari, Grunau, and Jin [29] mention that 4-coloring of a graph can be trivially used to compute a maximal independent set).
Advice in other models of computation with partial knowledge.
More generally, one can consider the impact of advice that captures some global knowledge about the whole input (or distribution of inputs) on the performance of algorithms. Similarly to the case of distributed computing, one can consider existential advice and advice that can be realized in a somewhat practical model of computation.
@@
Emek, Fraigniaud, Korman, and Rosén [15] introduced the notion of advice in the context of online algorithms. The authors studied the impact of advice in the competitive ratio of some classical online problems, namely metrical task systems and $k$-server, and they show tradeoffs between the number of bits of advice and the achievable competitive ratio. We refer to the survey by Boyar, Favrholdt, Kudahl, Larsen, and Mikkelsen [8] for more work on online algorithms with advice.
Mitzenmacher and Vassilvitskii [50] show that algorithms using advice realized by machine learning models can be used to provide algorithms that on average perform better than their traditional counterparts while keeping the same worst-case bounds. However, this particular advice model allows querying an oracle at each step. As such, in the context of online algorithms, it is in some sense both weaker than the existential advice for online algorithms (as it is produced by some machine learning model) and stronger as it gives a lot of bits of advice, and has access to the prefix of the input.
3 Preliminaries
@@
Let $G=(V,E)$ be a graph, where $V$ is the set of nodes and $E$ is the set of edges. We denote with $n$ the number of nodes of the graph, and with $\Delta$ its maximum degree. We may use the notation $V(G)$ to denote the set of nodes of $G$, and $E(G)$ to denote the set of edges of $G$. With $\operatorname{dist}_{G}(u,v)$ we denote the distance between $u\in V$ and $v\in V$ in $G$, that is, the length of the shortest path between $u$ and $v$ in $G$, where the length of a path is the number of edges of the path. For an integer $k$, the power graph $G^{k}$ of a graph $G$ is the graph that contains the same nodes as $G$, and there is an edge $\{u,v\}$ in $G^{k}$ if and only if $1\leq\operatorname{dist}_{G}(u,v)\leq k$. Two nodes $u$ and $v$ are neighbors in $G$ if there is an edge $\{u,v\}\in E(G)$.
@@
Lemma 3.1 (Lovász Local Lemma [56, 57, 16]).
@@ 3.1 LOCAL model
3.2 Locally checkable labelings
We formally define an LCL problem $\Pi$ as a tuple $(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},C,r)$ where each element of the tuple is defined as follows. The parameters $\Sigma_{\mathrm{in}}$ and $\Sigma_{\mathrm{out}}$ are finite sets of input and output labels, respectively. The parameter $r$ is a positive integer called checkability radius of $\Pi$, i.e., it determines how far in the graph each node needs to check in order to verify the validity of a given solution. The parameter $C$ determines the constraints of $\Pi$, that is, $C$ is a finite set of labeled graphs $H$ containing a vertex of eccentricity at most $r$, where each edge-endpoint pair $(uv,v)\in E_{H}\times V_{H}$ has a label $\ell_{\mathrm{in}}\in\Sigma_{\mathrm{in}}$ and a label $\ell_{\mathrm{out}}\in\Sigma_{\mathrm{out}}$.
Let $\Pi=(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},C,r)$ be an LCL problem and let $G=(V,E)$ be a graph where each edge-endpoint pair $(uv,v)$ is labeled with a label from $\Sigma_{\mathrm{in}}$. The task of solving $\Pi$ on $G$ requires labeling each edge-endpoint pair $(uv,v)\in E\times V$ with a label in $\Sigma_{\mathrm{out}}$ such that, for each node $v\in V$ it holds that the graph induced by the nodes at distance at most $r$ from $v$ and edges that have at least one endpoint at distance at most r1r-1 from $v$ is isomorphic to some (labeled) graph in $C$.
3.3 Advice schema
@@
Definition 3.2 (Advice Schema).
@@
  • There exists a LOCAL algorithm $\mathcal{A}$ that, for each $G\in\mathcal{G}$, if we label each node $v\in V$ with the bit-string $\ell_{G}(v)$, then $\mathcal{A}$ runs in at most $T$ rounds and outputs a valid solution for $\Pi$, where $T$ is a function of $\Delta$.
@@
  2. 2.
    If a subset of nodes receives bit-strings of the same length, and all the others receive bit-strings of length 0, then the schema is called subset fixed-length.
  3. 3.
    If a subset of nodes receives a bit-string of possibly different positive lengths, and all the others receive bit-strings of length 0, then the schema is called variable-length. Observe that an advice schema, as defined in Definition˜3.2 and without further assumptions, is variable-length.
@@
For uniform fixed-length advice schemas where all nodes receive one bit, we define the notion of sparsity, which captures the fact that the ratio between nodes receiving a $1$ and nodes receiving a 0 can be made arbitrarily small.
Definition 3.3 (Sparse schema).
@@
  • For any graph $G\in\mathcal{G}$, let $n_{0}$ be the number of nodes to which the schema (i.e., the function $f$ of Definition˜3.2) assigns a 0, and let $n_{1}$ be the number of nodes to which $f$ assigns a $1$. Then, $\frac{n_{1}}{n_{0}+n_{1}}\leq\varepsilon$.
@@
3.4 Composability
The main idea that we will use to devise our advice schemas is the following. Given some problem $\Pi$, it may be cumbersome to directly define an advice schema for it. Instead, it may be easier to do this operation gradually. In more detail, many problems can be solved as follows: first, find a solution for some subproblem; then, use the solution for the subproblem in order to solve the problem more easily. As a running example, consider the following problem $\Pi$.
@@ 3.4 Composability
  • It is required to output a coloring of the edges, say red and blue, such that each node has the same number of red and blue incident edges.
@@
  • Let $\Pi_{\mathrm{v}}$ be the problem of computing a $2$-coloring of the nodes, say, black and white.
  • Let $\Pi_{\mathrm{e}}$ be the problem of outputting a $2$-coloring of the edges, such that each node has the same number of red and blue incident edges, assuming that we are given as input a $2$-coloring of the nodes and a balanced orientation of the edges, and assuming that all nodes have even degree.
  • Let $\Pi_{\mathrm{o}}$ be the problem of orienting all edges such that each node has the same number of incoming and outgoing edges, assuming that all nodes have even degree.
Consider the following algorithm. First, solve $\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$. Then, solve $\Pi_{\mathrm{e}}$, by coloring red the edges oriented from black to white, and by coloring blue the edges oriented from white to black. Observe that this algorithm solves $\Pi$.
In other words, we decomposed $\Pi$ into three different subproblems: two of them are “hard” problems ($\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$), for which some advice is needed if we want to solve them fast, and one ($\Pi_{\mathrm{e}}$) is trivial once we are given as input the solution for the other two. The idea now is to devise two different advice schemas for $\Pi_{\mathrm{v}}$ and $\Pi_{\mathrm{o}}$ separately.
  • For $\Pi_{\mathrm{v}}$ there is a trivial advice schema: use $1$ bit to encode the $2$-coloring of the nodes.
  • For $\Pi_{\mathrm{o}}$ the advice schema is non-trivial, and we will see how it can be done in Appendix˜B by using $1$ bit per node.
@@
Once we have a composable schema for all subproblems that are used to solve our problem of interest, we apply, as a black box, a lemma (see Lemma˜E.1) to obtain a single variable-length schema for our problem. Then, again as a black box, we convert such a schema into a uniform fixed-length schema that uses a single bit per node (see Lemma˜E.3).
@@
  • We prove that such properties imply that we can compose many variable-length schemas into a single one that satisfies the same properties (see Lemma˜E.1);
  • We prove that the resulting variable-length schema implies a uniform fixed-length schema that uses a single bit per node (see Lemma˜E.3).
Note that the last two steps are done in a problem-independent way. The conditions that allow to combine multiple variable-length schemas are expressed in Definition˜3.4, that, on a high level, states the following. We want our variable-length schema to be tunable as a function of three parameters $\alpha$, $\gamma$, and $c$. The schema needs to satisfy that in each $\alpha$-radius neighborhood there are at most $\gamma$ bit-holding nodes, and that the number of bits held by these nodes is upper bounded by $c\alpha/\gamma^{3}$. Hence, a composable schema is a collection of schemas, one for each choice of parameters $\alpha$, $\gamma$, and $c$.
Definition 3.4 (Composability).
@@
4 Structural results and lower bounds
Our main goal here is to show that if we were able to solve all LCLs with some constant number of bits of advice per node, it would violate the Exponential-Time Hypothesis (ETH) [38]. However, to do that we need to first show a structural result, which is also of independent interest: algorithms that make use of advice can be made order-invariant, i.e., they do not need the numerical values of the identifiers.
4.1 Order-invariance
In general, in an advice schema our advice bits may depend on the numerical values of the identifiers, and the algorithm that solves the problem may make use of the advice bits and the numerical values of the identifiers. However, here we will show that we can essentially for free eliminate the dependency on the numerical values of the identifiers.
A distributed algorithm is called order-invariant [51] if the output remains the same if we change the numerical values of the identifiers but preserve their relative order. Put otherwise, an order-invariant algorithm may make use of the graph structure, local inputs (which in our case includes the advice), and the relative order of the identifiers, but not the numerical values of the identifiers.
Naor and Stockmeyer [51] showed that constant-time distributed algorithms that solve LCL problems can be made order-invariant, and since then order-invariant algorithms have played a key role in many lower-bound results related to constant-time distributed algorithms, see e.g. [34, 35].
Note that Theorem 4.1 below is applicable to LCL problems (they are component-wise defined, and in any LCL we have a bounded $\Delta$ and bounded alphabets $\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}}$), but also to many other problems that are not locally checkable. We emphasize that we will assume some bound on the maximum degree $\Delta$, but if we have an encoding schema that works for any $\Delta$, we can apply the following result separately for each $\Delta$.
Theorem 4.1.
Fix a maximum degree $\Delta$ and the number of advice bits $\beta$. Assume that $\Pi$ is a finite-input component-wise defined graph problem, in which the task is to label nodes with labels from some finite set $\Sigma_{\mathrm{out}}$. Assume that we can solve some graph problem $\Pi$ with some distributed algorithm in $T$ rounds using $\beta$ bits of advice. Then we can also solve $\Pi$ with an order-invariant algorithm in $T$ rounds using $\beta$ bits of advice.
Proof 4.2.
Let $E$ be the encoder that, given a graph $G$ (with some unique identifiers), produces $\beta$-bit advice strings that can then use to solve $\Pi$.
In the first steps, we follow the basic idea of Naor and Stockmeyer [51] to manipulate . Algorithm is a mapping from labeled radius-$T$ neighborhoods to local outputs; we write here $N^{T}(v)$ for the radius-$T$ neighborhood of node $v$, and we let $s=\Delta^{T+1}$ be an upper bound on the number of nodes in $N^{T}(v)$.
In general, we can encode all information in $N^{T}(v)$ as follows:
  1. 1.
    The set $X\subseteq\{1,2,\dotsc\}$ of unique identifiers present in the neighborhood, with $|X|\leq s$.
  2. 2.
So we can reinterpret as a function that maps a pair $(X,S)$ to the local output . But we can equally well interpret as a function $A$ that maps $X$ to a type $A(X)=F$, where $F$ is a function that maps $S$ to the local output $F(S)\in\Sigma_{\mathrm{out}}$; we simply let $A(X)(S)=$. The second interpretation turns out to be convenient.
Notice that we can always pad $X$ with additional identifiers (that use values larger than any value in $X$), so that we will have $|X|=s$.
Then we make the key observation: there are only finitely many different structures $S$. This follows from the assumptions that $\Delta$ is a constant, $\beta$ is a constant, $\Pi$ is a finite-input problem, and $T$ is a constant. Furthermore, $\Sigma_{\mathrm{out}}$ is finite, so $F$ is a mapping from a finite set to a finite set, and follows that there are only finitely many different types $F$. Let $m$ be the number of possible types, and identify the types with $1,2,\dots,m$.
Now let us stop for a moment and digest what we have learned so far: if our unique identifiers came from the set $I$, then we will have . That is, will then ignore the numerical values of the identifiers and only pay attention to the structure $S$. However, this is a big if; in general our identifiers can be arbitrary, and even adversarial.
Let us now continue; we will now modify the encoder $E$. We construct a new encoder $E^{\prime}$ that works as follows. Assume we are given a graph $G$, together with some assignment of unique identifiers. The encoder $E^{\prime}$ first constructs graph $G_{1}$ by renumbering the identifiers (but preserving their order) so that all identifiers in $G_{1}$ come from the monochromatic set $I$.
Now $H$ is a valid instance, and we can feed it to the encoder $E$, which will label it with $\beta$-bit labels so that if we apply to $H$ and these labels, we will correctly solve $\Pi$ in each component of $H$. In particular, we will correctly solve $\Pi$ in $G_{1}$. Moreover, in component $G_{1}$, algorithm will apply order-invariant algorithm $F^{*}$ in all neighborhoods, ignoring the numerical values of the identifiers.
If we now applied to solve $\Pi$ in $G$, it does not necessarily work. However, if we apply to solve $\Pi$ in $G$, it will behave in exactly the same way as applying in $G_{1}$. Hence, will also solve $\Pi$ correctly in $G$. Furthermore, is by construction order-invariant.
We note that the encoder $E^{\prime}$ constructed above is not practical or efficient (even if the original encoder $E$ is). As we will see next, merely knowing that exists can be sufficient.
4.2 Hardness assuming the Exponential-Time Hypothesis
In Appendix˜A we show that in graphs with sub-exponential growth, any LCL can be solved with one bit of advice per node. We now show that one bit does not suffice in general, assuming the Exponential-Time Hypothesis [38].
Recall that the Exponential-Time Hypothesis states that there is a positive constant $\delta>0$ such that 3-SAT cannot be solved in time $O(2^{\delta n})$, where $n$ is the number of variables in the 3-SAT instance.
Theorem 4.3.
Fix any $\beta$. Assume that for every LCL problem $\Pi$ there is some $T$ such that, on any input, $\Pi$ can be solved in $T$ rounds with $\beta$ bits of advice. Then the Exponential-Time Hypothesis is false.
Proof 4.4.
Consider some 3-SAT instance $\phi$, with $n$ variables and $m$ clauses. By applying the sparsification lemma by [39], we can assume w.l.o.g. that $m=O(n)$.
Now turn $\phi$ into a bipartite graph, with nodes representing variables on one side and nodes representing clauses on the other side. Each clause is connected by edges with the three variables it contains. Each node is labeled by its type (variable or clause) and each edge $vc$ is labeled to indicate whether the variable $v$ is negated in the clause $c$. This results in a bipartite graph with $n+m=O(n)$ nodes, and each clause-node has degree at most $3$. Hence, the total number of edges is at most $3m$.
Variable-nodes may have arbitrarily high degrees, but the sum of their degrees is bounded by $3m$. We replace each variable-node that has degree $d>3$ by a cycle of $d$ variable-nodes, with the new edges labeled by equality constraints. This results in a graph in which all nodes have degree at most $3$, and the total number of nodes is bounded by $3m+m=O(n)$. Let $G_{0}$ be the resulting graph, let $n_{0}=O(n)$ be the number of nodes in $G_{0}$, let $\Delta_{0}=3$ be the maximum degree, and let $\ell_{0}=O(1)$ be the number of node and edge labels that we used to encode the instance.
Now assume that we have defined an LCL problem $\Pi_{i}$ and a graph $G_{i}$ with $n_{i}$ nodes, maximum degree $\Delta_{i}$, and $\ell_{i}$ labels; the base case $i=0$ was presented above. Define a new LCL problem $\Pi_{i+1}$ and a new graph $G_{i+1}$ as follows. We contract edges in $G_{i}$ to construct $G_{i+1}$ so that we satisfy two properties: each node in $G_{i+1}$ represents $O(1)$ nodes of $G_{i}$, but the number of nodes is $n_{i+1}\leq n_{i}/2$. This can be achieved by e.g. greedily contracting edges, favoring the edges whose endpoints currently represent the smallest number of nodes of $G_{i}$.
We label the nodes of $G_{i+1}$ so that given the input labels of the nodes, we can also recover the original graph $G_{i}$. As each node represents a bounded number of original nodes, and the original graph had maximum degree $\Delta_{i}$, the new graph will have maximum degree $\Delta_{i+1}$ that only depends on $i$. Also, the number of labels $\ell_{i+1}$ will be bounded by a constant that only depends on $i$. To see that everything can be indeed encoded with constant-size labels, note that each new node $v$ in $G_{i+1}$ can have its own local numbering of the original nodes $v^{\prime}$ that were contracted to $v$, that is, each node $v^{\prime}$ can be represented as a pair $(v,a)$, where $v$ is a new node and $a$ is a sequence number. This way, the new label of node $v$ can use constant-size triples of the form $(a,b,x)$ to encode that graph $G_{i}$ had an edge from $(v,a)$ to $(v,b)$ with label $x$, and the new edge label of edge $(u,v)$ can also use constant-size triples of the form $(a,b,x)$ to encode that graph $G_{i}$ had an edge from $(u,a)$ to $(v,b)$ with label $x$. Finally, the new node label of $v$ will use constant-size pairs of the form $(a,x)$ to encode that the node label of $(v,a)$ in $G_{i}$ was $x$. This way we can use constant-size node and edge labels in $G_{i+1}$ to encode the structure and node and edge labels of $G_{i}$.
By assumption, there exists a distributed algorithm that solves $\Pi_{i}$ with $\beta$ bits of advice per node. Using Theorem˜4.1, we can also assume that is order-invariant. Hence, is a finite function, and we can compute in constant time (where the constant depends on $i$, which depends on $\beta$ and $\varepsilon$, but is independent of $n$).
Now, we simply try out all possible strings of advice; there are at most $2^{\varepsilon n}$ such strings. For each advice combination, we try to apply to solve $\Pi_{i}$ in $G_{i}$; we simulate at each of the $n_{i}=O(n)$ nodes. If and only if $\phi$ is satisfiable, we will find an advice string such that succeeds in solving $\Pi_{i}$. The overall running time is $O(2^{\varepsilon n}\cdot n\cdot f(\varepsilon,\beta))$.
Appendix A LCLs on graphs with sub-exponential growth